OpenStudy (anonymous):

if log1/2 X <3 and log4 X<2, then (a) 1/8

OpenStudy (anonymous):

see.. in this question a little thing to take care of is base.. you increase the power of a number and u get a larger no... but that is not the case with numbers between 0 & 1.. so if u increase the power of 1/2 & u will be getting lesser no. than previous.. so we go like this.. suppose log1/2X=3 from here u can calculate X=(1/2)^3=1/8 now what you want is log1/2X should be lesser than 3 hence what..? X should be greater than 1/8.. that is the point.. now from 2nd expression you need log4X<2 hence what u need is X<16.. now the final solution should be intersection of both..hence the 1st option is correct.. remember the following graph & u will never confuse ever..